Given f(x)=x^3+kx+2, and x+1 is a factor of f(x), then what is the value of k?
Found 2 solutions by Edwin McCravy, ikleyn:
Answer by Edwin McCravy(20056)

Instead of doing your homework for you, I'll do one exactly like yours step-by-step
so you can use it as a model to do yours by.
Given f(x)=x^3+kx-36, and x-3 is a factor
of f(x), then what is the value of k?
We divide f(x) by x-3, using synthetic division (be sure to change the sign of
-3 to +3 when using synthetic division).

 3 | 1   0   k   -36
   |     3   9  3k+27
     1   3  k+9 3k-9

The remainder 3k-9 must be 0, so 3k-9 = 0
                                   3k = 9
                                    k = 3

Edwin

Answer by ikleyn(52788)

.
Given f(x)=x^3+kx+2, and x+1 is a factor of f(x), then what is the value of k?
~~~~~~~~~~~~~~~~~

The fact that (x+1) is a factor of the polynomial f(x) = x^3 + kx + 2, means that the value of -1
is the root of this polynomial: f(-1) = 0,  due to the Remainder theorem.


So, we substitute x= -1 into the polynomial and equate it to zero


    (-1)^3 + k*(-1) + 2 = 0.


It is an equation to determine the value of k.  So, simplify and find k


    -1 -k + 2 = 0

    -1 + 2 = k

     k = 1.


ANSWER.  k = 1.

Solved.

-------------

If you want to see many other similar  (and different)  solved problems,  look into the lesson
    - Finding unknown coefficients of a polynomial having given info about its polynomial divisors
in this site.

On the Remainder theorem, see the lessons
    - Divisibility of polynomial f(x) by binomial (x-a) and the Remainder theorem
    - Typical problems on the Remainder theorem
    - Advanced problems on the Remainder theorem
    - Finding unknown coefficients of a polynomial having given info about its polynomial divisors
    - Finding unknown coefficients of a polynomial based on some given info about its roots
in this site.